OpenStudy (cherie_magee):

Given: x + y = 15. Which of the following are ordered pairs of this equation? {(-3, 14),(0, 15),(3, 16)} {(-3, 16),(0, 15),(3, 14)} {(16, -3),(15, 0),(14, 3)}

terenzreignz (terenzreignz):

Well, just pick the set containing pairs that when you add up, equal 15.

terenzreignz (terenzreignz):

Consider the first one... it's not your answer, because it has the pair (-3, 14) which, when you add up... \[\large -3 + 14 = 11 \ne 15\]

OpenStudy (cherie_magee):

in that case none of these are right @terenzreignz

terenzreignz (terenzreignz):

Are these sets of three pairs?

terenzreignz (terenzreignz):

or 9 individual pairs?

OpenStudy (cherie_magee):

sets of three pairs @terenzreignz

terenzreignz (terenzreignz):

Then in that case, none of these would be right...

OpenStudy (cherie_magee):

i forgot to put the 1/3 infront of x my bad @terenzreignz

terenzreignz (terenzreignz):

<eye twitching>

OpenStudy (cherie_magee):

lol just answer the question kml @terenzreignz

terenzreignz (terenzreignz):

Well, in that case, what you do is just plug in, instead of adding... \[\Large \frac13x + y = 15\]

terenzreignz (terenzreignz):

For instance, that first set again... is not the solution, because when we plug in (-3 , 14) x = -3 y = 14 \[\Large \frac13(-3)+14 = -1 + 14 = 13 \ne15\]

OpenStudy (cherie_magee):

um i dont know how to do that yeah mhm so tell me @terenzreignz

terenzreignz (terenzreignz):

Okay, again... step-by-step... let's test out the first point of the first set of three points... That would be (-3, 14) So x = -3 and y = 14 Let's test it out on \[\Large \frac13\color{red}x+\color{blue}y=15\]

terenzreignz (terenzreignz):

\[\Large \frac13(\color{red}{-3})+\color{blue}{14}\]\[\Large = -1+14\]\[\Large = 13\]\[\Large \ne 15\] We need it to be 15, but this pair doesn't fit. Hence, the first set isn't the answer :P

OpenStudy (cherie_magee):

oooook i get it now thank ya sweetie @terenzreignz

terenzreignz (terenzreignz):

It's 'Terence', sweetie :P And no problem, now tell me your answer :3

OpenStudy (cherie_magee):

ok smarty but i ended up with with the last set @terenzreignz

terenzreignz (terenzreignz):

It's Terence -.- And really now? I fail to see how even the pair (16 , -3) fits into all of this...

terenzreignz (terenzreignz):

The pair (16, -3) is in the last set, no? x = 16 y = -3 \[\Large \frac13\color{red}x+\color{blue}y=15\] \[\Large \frac13(\color{red}{16})\color{blue}{-3}=15\]\[\Large \color{green}{\frac{7}{3}}=15\]

terenzreignz (terenzreignz):

clearly not true... *sweetie* :P

OpenStudy (cherie_magee):

never mind its the second one i freaked up and put x and y in parentheses an did that first @terenzreignz

terenzreignz (terenzreignz):

Okay, well, since I already said the first and last were wrong, it has to be the second one, *sweetie* Good job anyway :P

OpenStudy (cherie_magee):

i wasn't even looking at that TERENCE @terenzreignz

terenzreignz (terenzreignz):

U mad? LOL Peace ^_^

OpenStudy (cherie_magee):

k bye kml @terenzreignz

OpenStudy (cherie_magee):

this ended up being wrong TERENCE @terenzreigns

terenzreignz (terenzreignz):

Impossible, unless there was something you forgot to put up again...

OpenStudy (cherie_magee):

thats what i thought so i worked it out step by step and i still got the same answer these people must be dumb or its something i'm missing @terenzreignz

terenzreignz (terenzreignz):

Interestingly, what was apparently the correct answer?